# InfiJanice: Joint Analysis and In-situ Correction Engine for Quantization-Induced Math Degradation in Large Language Models

### v1

@@ 4.4 Ablation Study
To isolate the contributions of each component in our quantization recovery pipeline, we perform a series of ablation studies. Specifically, we investigate the impact of (1) using subsets focused on a single error type, (2) sampling problems at random rather than targeting those where the quantized model fails, and (3) training with non-step-wise generated data by directly adopting full-precision model solutions as positive cases. We evaluate how each variation affects the model’s ability to recover reasoning performance on mathematical benchmarks.
@@ 4.5 Discussion
Drawing on the findings in Sections 4.3 and 4.4, together with the trends visualized in Figure 2, we highlight three key takeaways:
Targeted recovery with our "Silver Bullet" datasets. Fine-tuning on our curated "Silver Bullet" split restores the model’s accuracy on MATH while simultaneously boosting performance on the less demanding GSM8K benchmark. Crucially, this intervention leaves broader reasoning skills, as measured by MMLU, unchanged.
Quantization erodes Execution & Reasoning skills in mathematics. Weight–activation quantization chiefly harms the model’s ability to carry out intermediate calculations and boundary-condition reasoning. These deficits propagate: a single arithmetic slip or mis-handled constraint can invalidate an otherwise correct derivation, explaining many of the failures observed on math-focused tasks.
Step-wise positives outperform naive alternatives. Ablation studies show that regenerating the solution sequence from the first erroneous step onward and using that trace as a positive example yields larger gains than either (i) simply adopting the full-precision model’s complete derivation or (ii) randomly mixing correct and incorrect traces from the quantized model. Focusing supervision precisely where the model falters is therefore the most effective strategy for capability recovery.
@@ 5 Conclusion
In this study, we present a systematic investigation into how quantization impacts the mathematical reasoning capabilities of LLMs across varying model scales and task complexities. We demonstrate that aggressive compression significantly compromises performance on benchmarks such as GSM8K, MATH, and AIME. To address this degradation, we introduce a format alignment protocol that regularizes model outputs into structured reasoning traces, along with a scalable assessment pipeline that categorizes fine-grained error types with high reliability. Building upon this analysis, we curate high-impact "Silver Bullet" datasets targeting model-specific weaknesses. Fine-tuning quantized models via Direct Preference Optimization (DPO) on these datasets allows us to recover performance using only minimal compute resources. Our approach restores accuracy while preserving the efficiency gains of quantization. Overall, our findings highlight a practical path for deploying quantized LLMs without sacrificing core reasoning abilities. Future work will extend our methodology to domains such as code generation, tool usage, and symbolic reasoning, while further automating the pipeline for broader and more scalable applications.
@@ A.1 AWQ
AWQ (Activation–Aware Weight Quantization) compensates for the long-tailed distribution of activations before the weight tensor is discretised. Let $\mathbf{A}\!\in\!\mathbb{R}^{B\times d}$ be the mini-batch activations and $\mathbf{W}\!\in\!\mathbb{R}^{d\times m}$ the corresponding weights. A positive scale vector $\boldsymbol{\gamma}\!\in\!\mathbb{R}^{d}_{+}$ is chosen such that
$\widetilde{\mathbf{Y}}=\bigl{(}\mathbf{A}\odot\boldsymbol{\gamma}^{-1}\bigr{)}\bigl{(}\boldsymbol{\gamma}\odot Q(\mathbf{W})\bigr{)}^{\top},\quad\gamma_{k}=\bigl{(}\operatorname{mean}|\mathbf{A}_{\!,k}|^{\,\alpha}\bigr{)}\bigl{(}\operatorname{mean}|\mathbf{W}_{k,\!:\!}|^{-\beta}\bigr{)},$
where $(\alpha,\beta)\!\in\![0,1]$ control the balance between activation and weight magnitudes and $Q(\cdot)$ denotes an asymmetric 4-bit quantiser. Because the rescaling is folded back into $\mathbf{W}$, the forward pass is identical to the unscaled INT4 kernel and incurs no extra latency.
@@ A.2 GPTQ
GPTQ formulates post-training quantisation as a blockwise least-squares problem over a small calibration set $\mathcal{C}=\{\mathbf{A}^{(i)}\}_{i=1}^{|\mathcal{C}|}$:
@@
where $\mathcal{Q}$ is the set of weight tensors representable by the target bit-width. The optimisation proceeds greedily over 128-channel blocks. After quantising one block, GPTQ updates the remaining full-precision weights with a rank-$r$ approximation of the corresponding Hessian inverse, cheaply computed from second-order activation statistics. This strategy yields near-optimal INT4 weights with negligible calibration cost.
@@ A.3 SmoothQuant
The scales $\{\sigma_{k}\}$ are obtained by minimising the worst-case per-channel quantisation error across the calibration set, typically using a few thousand tokens. [^1] We use 2 04820482\,0482 048 tokens, following Xiao et al. [2023]. Once trained, the scales are fused into $\mathbf{W}$ and the model runs on standard INT8 kernels without auxiliary tensors or runtime re-scaling.
@@
All three methods adopt per-channel affine quantisation. AWQ and GPTQ target 4-bit weights and retain FP16 activations, whereas SmoothQuant yields a fully INT8 model. We keep the original hyper-parameters recommended by the respective authors to ensure reproducibility across codebases.
@@ Appendix B Prompt
Problem Statement (from PRM-800K Dataset Training Subsets) The product of the proper positive integer factors of $n$ can be written as $n^{(ax+b)/c}$, where $x$ is the number of positive divisors $n$ has, $c$ is a positive integer, and the greatest common factor of the three integers $a$, $b$, and $c$ is $1$. What is $a+b+c$? [Gold Answer: 1]
@@ Appendix C Human Annotation Guidebook
This guideline specifies the manual verification protocol applied to disagreement cases that survive the automated evaluation pipeline—namely the expert-LLM judges and the five-model majority vote. Annotators produce the final ground-truth verdict (error type, error step, explanation, confidence) for every instance in which
@@
  • answers.json: correct answer, full-precision answer, quantised answer.
@@
    Locate first divergence. Read fp_trace and qt_trace in parallel and find the earliest step where the quantised trace deviates from valid reasoning.
@@
  • Project leads re-annotate 2 % of “agree” cases and all needs_second_opinion cases; overall agreement $<95\%$ triggers targeted review.
@@ D.1 Example of Execution Errors
In Case #93, the task was to find the value of $c$ for which the circle defined by $x^{2}-10x+y^{2}+6y+c=0$ has a radius of $1$. The correct approach involves completing the square, resulting in $(x-5)^{2}+(y+3)^{2}=-c+34$, and solving $-c+34=1$ to find $c=33$ However, the GPTQModel made a computational error by incorrectly rearranging the equation as $134-c=$ 1 leading to the wrong answer $c=35$. This error stemmed from mishandling the sign of $c$ during algebraic manipulation, highlighting the importance of careful sign management in equation solving. The case #93 shows the Execution Errors.
Problem #93 For what value of $c$ will the circle with equation $x^{2}-10x+y^{2}+6y+c=0$ have a radius of length $1$? [Gold Answer: 33]
Response of LLaMA-3.1-8B-Intruct-BF16 (Vanilla) 1. To find the radius of a circle, I need to complete the square for both $x$ and $y$ terms in the equation. 2. I can rewrite the equation as $(x^{2}-10x)+(y^{2}+6y)=-c$. 3. To complete the square for $x^{2}-10x$, I add $(10/2)^{2}=25$ to both sides. 4. To complete the square for $y^{2}+6y$, I add $(6/2)^{2}=9$ to both sides. 5. This gives $(x^{2}-10x+25)+(y^{2}+6y+9)=-c+25+9$. 6. I factor the left side as $(x-5)^{2}+(y+3)^{2}$. 7. The right side simplifies to $-c+34$. 8. Therefore the equation becomes $(x-5)^{2}+(y+3)^{2}=-c+34$. 9. The radius of the circle is the square root of the right-hand side. 10. Setting the radius to $1$ gives $-c+34=1^{2}=1$. 11. Solving for $c$ yields $c=34-1=33$. [Answer: 33]
Response of LLaMA-3.1-8B-Intruct-W4A16 (GPTQ Model) 1. To find the radius, I first complete the square for the $x$ and $y$ terms and rewrite the equation in the form $(x-h)^{2}+(y-k)^{2}=r^{2}$, where $(h,k)$ is the centre and $r$ is the radius. 2. I begin with the $x$ terms, adding and subtracting the same value to form a perfect square. 3. Because the coefficient of $x$ is $-10$, I add and subtract $\bigl{(}-10/2\bigr{)}^{2}=25$ to obtain $x^{2}-10x+25-25$. 4. Factoring gives $(x-5)^{2}$ and the equation becomes $(x-5)^{2}-25+y^{2}+6y+c=0$. 5. I repeat the process for the $y$ terms, adding and subtracting $\bigl{(}-6/2\bigr{)}^{2}=9$ to get $y^{2}+6y+9-9$. 6. Factoring yields $(y+3)^{2}$ and the equation is now $(x-5)^{2}-25+(y+3)^{2}-9+c=0$. 7. Combining constants, I have $(x-5)^{2}+(y+3)^{2}-34+c=0$. 8. Setting $r^{2}=1$ requires $-34+c=1$; solving gives $c=35$. 9. Hence $c=35$. [Answer: 35]
@@ D.3 Example of No Error
[…]
@@
Response of LLaMA-3.1-8B-Intruct-W4A16 (GPTQ Model) 1. To find the probability of rolling a prime number on a fair 6-sided die, I need to know how many prime numbers appear on such a die and how many faces there are in total. 2. A prime number is a positive integer divisible only by 1 and itself. The prime numbers on a 6-sided die are 2, 3,232,\,3,2 , 3 , and $5$. 3. There are 3 prime numbers on a 6-sided die, so the probability of rolling a prime number is $3$ out of $6$, or $1/2$. 4. Converting this fraction to a decimal, I divide $1$ by $2$ to obtain $0.5$. 5. Alternatively, multiplying $1$ by $100$ and dividing by $2$ gives $50\%$. 6. Therefore, the probability of rolling a prime number on a fair 6-sided die is $0.5$ or $50\%$. [Answer: 0.5]
Judgement of Deepseek-R1 (Primary Expert Model) The quantized model’s answer ($0.5$) is numerically equivalent to the correct answer ($1/2$). All reasoning steps are logically valid and mathematically correct, with no conceptual, procedural, or computational errors detected. [No_Error: [][][ ]]
@@ D.4 Example of Conflicting Judgments
Response of LLaMA-3.1-8B-Intruct-W4A16 (GPTQ Model) 1. To solve this inequality, I need to isolate $x$ on one side. 2. I can start by multiplying both sides by $4$, since $4$ is a common denominator for the fractions. 3. This gives me $1<\frac{4x}{7}<\frac{4}{3}$. 4. Now I can multiply both sides by $7$, since $7$ is a common denominator for the fractions. 5. This gives me $7<4x<\frac{28}{3}$. 6. To make the inequality easier to read, I can divide both sides by $4$. 7. This gives me $\frac{7}{4}<x<\frac{7}{1}$. 8. I can simplify the fractions by writing them as mixed numbers or decimals. 9. This gives me $71.75<x<$ 7. 10. Since $x$ is an integer, I need to find the integers that satisfy this inequality. 11. The smallest integer greater than $1.75$ is $2$. 12. The largest integer less than $7$ is $6$. 13. Therefore, the integers that satisfy the inequality are $2,3,4,5,6$. [Answer: 2,3,4,5,6]
Judgement of Deepseek-R1 (Primary Expert Model) When multiplying $4/3$ by $7$, the correct result is $28/3\approx 9.33$. In step 5 the quantized model treats $7\times(4/3)$ as $28/3$ but, in step 7, fails to divide this by $4$ correctly, giving an upper bound of $7$ instead of $7/3\approx 2.33$. [Execution_Errors: computational_error]
Judgement of Deepseek-V3 The quantized model incorrectly multiplies both sides by $4$ instead of $7$, the proper common denominator. This leads to a procedural error that propagates through the subsequent steps. [Method_Errors: procedural_error]
@@
Judgement of Qwen-Max The model simplifies $\tfrac{7}{4}<x<\tfrac{28}{3}$ to $71.75<x<$ 7. The upper bound $\tfrac{28}{3}$ is approximately $9.33$, not $7$. [Conceptual_Errors: conceptual_misunderstanding]
@@ Appendix F Capability Restoration Results
Appendix G Limitations and Future Works
Our analysis is restricted to the MATH benchmark and to models that fit within our available computational budget. These constraints preclude a full ablation study, limit the breadth of case analyses, and leave the behaviour of larger-scale quantized models unexplored. Consequently, some observations—such as occasional gains on specific sub-steps of mathematical proofs—could not be investigated in depth.
In the future, we intend to refine our error diagnostics, focus on the most fragile capabilities, and incorporate expert annotation to enrich qualitative insights. Over a longer horizon we will expand the evaluation to additional reasoning domains—such as code generation and tool use—while systematically comparing quantized and full-precision models across parameter scales. Together, these efforts should yield a more complete understanding of how low-bit representations shape reasoning behavior and how their weaknesses can be mitigated.

### v3

@@ 4.4 Ablation Study
To isolate the contributions of each component in our quantization recovery pipeline, we perform a series of ablation studies. Unless otherwise noted, all runs fix the training budget, optimizer, prompts, and decoding policy. We compare four variants of our training pairs (the failure subset refers to instances the quantized model answers incorrectly under baseline evaluation):
  • All-Step: all error cases from the failure subset; step-aligned supervision resumes at the first-error step (i.e., the first step where the model deviates from the gold solution) and continues step-wise along the gold trajectory to the final answer.
  • CE/ME/EE-Step: identical to All Errors-Step but restricted to a single error type (Conceptual-Error / Method-Error / Execution-Error), with supervision resuming from the first-error step.
  • Random-NonStep: size-matched random sampling from the math corpus, independent of whether the quantized model fails; positives are the gold solutions, and no first-error resuming is applied.
  • ALL-NonStep: all error cases from the failure subset, but without resuming from the first-error step; positives are the full gold solutions.
@@ 4.5 Discussion
Synthesizing the results from Sections 4.3 and 4.4 together with the trends in Figure 2 and Table 2, we draw three main conclusions:
(i) Targeted recovery with our "Silver Bullet" datasets.
Fine-tuning on the compact failure-targeted split substantially restores performance on MATH500 while also boosting GSM8K, and does so without hurting broad-domain reasoning as measured by MMLU. This confirms that a small, carefully constructed dataset can achieve capability recovery rather than memorization, aligning with our design goal of lightweight in-situ correction.
(ii) Quantization disproportionately erodes procedural and executional skills.
Our error taxonomy shows that weight–activation quantization mainly increases method and execution errors—such as carrying out multi-step arithmetic or handling boundary conditions—rather than high-level conceptual reasoning. Because these mistakes often occur early, they propagate to invalidate otherwise correct derivations, explaining the steep drop on math-centric tasks.
(iii) Step-wise positives outperform naive alternatives.
Ablation results (Table 2) demonstrate that regenerating solutions from the first erroneous step onward yields larger gains than either (i) adopting the full-precision model’s complete derivation (All-NS), or (ii) training on size-matched random problems (Rand-NS). Focusing supervision precisely at the point of failure proves to be the most effective strategy for restoring mathematical reasoning.
@@ 5 Conclusion
In this study, we present a systematic study of quantization-induced degradation in the mathematical reasoning abilities of large language models, revealing that low-bit post-training quantization especially harms smaller models’ procedural and execution skills. To address this, we propose a lightweight recovery pipeline that combines step-aligned error analysis with targeted fine-tuning on compact, automatically constructed “Silver Bullet” datasets. Experiments show that, with minimal data and compute, quantized models can recover reasoning performance to match their full-precision counterparts while preserving efficiency and general capabilities. Our approach offers a practical and extensible solution for deploying quantized LLMs in resource-constrained settings, and opens avenues for robust reasoning restoration in broader domains.
@@ A.1 AWQ
AWQ (Activation–Aware Weight Quantization) compensates for the long‐tailed distribution of activations before the weight tensor is discretised. Let $\mathbf{A}\!\in\!\mathbb{R}^{B\times d}$ be the mini‑batch activations and $\mathbf{W}\!\in\!\mathbb{R}^{d\times m}$ the corresponding weights. A positive scale vector $\bm{\gamma}\!\in\!\mathbb{R}^{d}_{+}$ is chosen such that
$\widetilde{\mathbf{Y}}=\bigl(\mathbf{A}\odot\bm{\gamma}^{-1}\bigr)\bigl(\bm{\gamma}\odot Q(\mathbf{W})\bigr)^{\top},\quad\gamma_{k}=\bigl(\operatorname{mean}|\mathbf{A}_{\!,k}|^{\,\alpha}\bigr)\bigl(\operatorname{mean}|\mathbf{W}_{k,\!:\!}|^{-\beta}\bigr),$
where $(\alpha,\beta)\!\in\![0,1]$ control the balance between activation and weight magnitudes and $Q(\cdot)$ denotes an asymmetric 4‑bit quantiser. Because the rescaling is folded back into $\mathbf{W}$, the forward pass is identical to the unscaled INT4 kernel and incurs no extra latency.
@@ A.2 GPTQ
GPTQ formulates post‑training quantisation as a blockwise least‑squares problem over a small calibration set $\mathcal{C}=\{\mathbf{A}^{(i)}\}_{i=1}^{|\mathcal{C}|}$:
@@
where $\mathcal{Q}$ is the set of weight tensors representable by the target bit‑width. The optimisation proceeds greedily over 128‑channel blocks. After quantising one block, GPTQ updates the remaining full‑precision weights with a rank‑$r$ approximation of the corresponding Hessian inverse, cheaply computed from second‑order activation statistics. This strategy yields near‑optimal INT4 weights with negligible calibration cost.
@@ A.3 SmoothQuant
The scales $\{\sigma_{k}\}$ are obtained by minimising the worst‑case per‑channel quantisation error across the calibration set, typically using a few thousand tokens. [^1] We use 2 0482\,048 tokens, following Xiao et al. (2023). Once trained, the scales are fused into $\mathbf{W}$ and the model runs on standard INT8 kernels without auxiliary tensors or runtime re‑scaling.
@@
All three methods adopt per‑channel affine quantisation. AWQ and GPTQ target 4‑bit weights and retain FP16 activations, whereas SmoothQuant yields a fully INT8 model. We keep the original hyper‑parameters recommended by the respective authors to ensure reproducibility across codebases.
@@ Appendix B Prompt
Problem Statement (from PRM-800K Dataset Training Subsets) The product of the proper positive integer factors of $n$ can be written as $n^{(ax+b)/c}$, where $x$ is the number of positive divisors $n$ has, $c$ is a positive integer, and the greatest common factor of the three integers $a$, $b$, and $c$ is 11. What is $a+b+c$? [Gold Answer: 1]
@@ Appendix C Human Annotation Guidebook
This guideline specifies the manual verification protocol applied to disagreement cases that survive the automated evaluation pipeline—namely the expert‐LLM judges and the five-model majority vote. Annotators produce the final ground-truth verdict (error type, error step, explanation, confidence) for every instance in which
@@
  • answers.json: correct answer, full-precision answer, quantized answer.
@@
    Locate first divergence. Read fp_trace and qt_trace in parallel and find the earliest step where the quantized trace deviates from valid reasoning.
@@
  • Project leads re-annotate 2 % of “agree” cases and all needs_second_opinion cases; overall agreement $95\%$ triggers targeted review.
@@ D.1 Example of Execution Errors
In Case #93, the task was to find the value of $c$ for which the circle defined by $x^{2}-10x+y^{2}+6y+c=0$ has a radius of $1$. The correct approach involves completing the square, resulting in $(x-5)^{2}+(y+3)^{2}=-c+34$, and solving c+34=1-c+34=1 to find $c=33$ . However, the GPTQModel made a computational error by incorrectly rearranging the equation as 34c=134-c=1 leading to the wrong answer $c=35$. This error stemmed from mishandling the sign of $c$ during algebraic manipulation, highlighting the importance of careful sign management in equation solving. The case #93 shows the Execution Errors.
Problem #93 For what value of $c$ will the circle with equation $x^{2}-10x+y^{2}+6y+c=0$ have a radius of length 11? [Gold Answer: 33]
Response of LLaMA-3.1-8B-Intruct-BF16 (Vanilla) 1. To find the radius of a circle, I need to complete the square for both $x$ and $y$ terms in the equation. 2. I can rewrite the equation as $(x^{2}-10x)+(y^{2}+6y)=-c$. 3. To complete the square for $x^{2}-10x$, I add $(10/2)^{2}=25$ to both sides. 4. To complete the square for $y^{2}+6y$, I add $(6/2)^{2}=9$ to both sides. 5. This gives $(x^{2}-10x+25)+(y^{2}+6y+9)=-c+25+9$. 6. I factor the left side as $(x-5)^{2}+(y+3)^{2}$. 7. The right side simplifies to c+34-c+34. 8. Therefore the equation becomes $(x-5)^{2}+(y+3)^{2}=-c+34$. 9. The radius of the circle is the square root of the right-hand side. 10. Setting the radius to 11 gives $c+34=1^{2}=1$. 11. Solving for $c$ yields c=341=33c=34-1=33. [Answer: 33]
Response of LLaMA-3.1-8B-Intruct-W4A16 (GPTQ Model) 1. To find the radius, I first complete the square for the $x$ and $y$ terms and rewrite the equation in the form $(x-h)^{2}+(y-k)^{2}=r^{2}$, where $(h,k)$ is the centre and $r$ is the radius. 2. I begin with the $x$ terms, adding and subtracting the same value to form a perfect square. 3. Because the coefficient of $x$ is 10-10, I add and subtract $\bigl(-10/2\bigr)^{2}=25$ to obtain $x^{2}-10x+25-25$. 4. Factoring gives $(x-5)^{2}$ and the equation becomes $(x-5)^{2}-25+y^{2}+6y+c=0$. 5. I repeat the process for the $y$ terms, adding and subtracting $\bigl(-6/2\bigr)^{2}=9$ to get $y^{2}+6y+9-9$. 6. Factoring yields $(y+3)^{2}$ and the equation is now $(x-5)^{2}-25+(y+3)^{2}-9+c=0$. 7. Combining constants, I have $(x-5)^{2}+(y+3)^{2}-34+c=0$. 8. Setting $r^{2}=1$ requires 34+c=1-34+c=1; solving gives $c=35$. 9. Hence $c=35$. [Answer: 35]
@@ D.3 Example of No Error
Most ambiguities stem from multiple notations for the same numeric value—particularly decimals versus fractions and natural versus programming syntax—e.g., $1/2$, $0.5$, $\frac${1}{2}, $5\mathrm{E}{-01}$, and $5\times 10^{-1}$.
@@
Response of LLaMA-3.1-8B-Intruct-W4A16 (GPTQ Model) 1. To find the probability of rolling a prime number on a fair 6-sided die, I need to know how many prime numbers appear on such a die and how many faces there are in total. 2. A prime number is a positive integer divisible only by 1 and itself. The prime numbers on a 6-sided die are 2, 3,2,\,3, and $5$. 3. There are 3 prime numbers on a 6-sided die, so the probability of rolling a prime number is 33 out of 66, or $1/2$. 4. Converting this fraction to a decimal, I divide 11 by 22 to obtain $0.5$. 5. Alternatively, multiplying 11 by $100$ and dividing by 22 gives $50\%$. 6. Therefore, the probability of rolling a prime number on a fair 6-sided die is $0.5$ or $50\%$. [Answer: 0.5]
Judgement of Deepseek-R1 (Primary Expert Model) The quantized model’s answer ($0.5$) is numerically equivalent to the correct answer ($1/2$). All reasoning steps are logically valid and mathematically correct, with no conceptual, procedural, or computational errors detected. [No_Error: $[]$]
@@ D.4 Example of Conflicting Judgments
Response of LLaMA-3.1-8B-Intruct-W4A16 (GPTQ Model) 1. To solve this inequality, I need to isolate $x$ on one side. 2. I can start by multiplying both sides by 44, since 44 is a common denominator for the fractions. 3. This gives me $1<\frac{4x}{7}<\frac{4}{3}$. 4. Now I can multiply both sides by 77, since 77 is a common denominator for the fractions. 5. This gives me $7<4x<\frac{28}{3}$. 6. To make the inequality easier to read, I can divide both sides by $4$. 7. This gives me $\frac{7}{4}<x<\frac{7}{1}$. 8. I can simplify the fractions by writing them as mixed numbers or decimals. 9. This gives me $1.75<x<7$. 10. Since $x$ is an integer, I need to find the integers that satisfy this inequality. 11. The smallest integer greater than $1.75$ is 22. 12. The largest integer less than 77 is 66. 13. Therefore, the integers that satisfy the inequality are $2,3,4,5,6$. [Answer: 2,3,4,5,6]
Judgement of Deepseek-R1 (Primary Expert Model) When multiplying $4/3$ by 77, the correct result is $28/3\approx 9.33$. In step 5 the quantized model treats $7\times(4/3)$ as $28/3$ but, in step 7, fails to divide this by 44 correctly, giving an upper bound of 77 instead of $7/3\approx 2.33$. [Execution_Errors: computational_error]
Judgement of Deepseek-V3 The quantized model incorrectly multiplies both sides by 44 instead of 77, the proper common denominator. This leads to a procedural error that propagates through the subsequent steps. [Method_Errors: procedural_error]
@@
Judgement of Qwen-Max The model simplifies $\tfrac{7}{4}<x<\tfrac{28}{3}$ to $1.75<x<7$. The upper bound $\tfrac{28}{3}$ is approximately $9.33$, not 77. [Conceptual_Errors: conceptual_misunderstanding]
@@ Appendix F Capability Restoration Results
Appendix G The Usage of LLM
In this work, Large Language Models (LLMs) were used as auxiliary tools to support our research process, but not to generate novel scientific content. Specifically, their usage includes:
  • Editing and polishing. LLMs were employed for minor grammar checking, improving clarity, and rephrasing sentences for readability in the manuscript. All scientific ideas, methodology, and experiments were designed and written by the authors.
  • Facilitating annotation. During the construction of our automated error-assessment pipeline, LLMs were used as expert judges to classify error types in reasoning traces. Their outputs were combined via majority voting and, when necessary, verified by human annotators to ensure reliability.
  • Experiment assistance. LLMs were queried to simulate baseline reasoning traces for building our contrastive “Silver Bullet” datasets, which were later curated, filtered, and validated by the authors. This step complements human effort by accelerating the generation of positive examples.
We emphasize that all key contributions—including research ideas, methodology design, experimental execution, and analysis—were conceived and implemented by the authors.